OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. 3x + 4y = -2 and 2x - 3y = 10

OpenStudy (jhannybean):

3x + 4y = -2 2x - 3y = 10 --------------- See if you can get rid of one variable first. Lets start with y. 3(3x + 4y = -2) 4(2x - 3y = 10) ----------------- 9x + 12y = -6 8x - 12y = 40 ----------------- 17x + 0y = 34 17x = 34 x = 34/17 = 2 now plug in this value for y and you will have your two variables.

OpenStudy (jhannybean):

Plugging in for y take any of the two equations and use x=2 in place of x. 2x - 3y = 10 x=2 2(2) -3y = 10 4 - 3y = 10 -4 +4 -3y = 10 - 4 -3y = 6 y=??

OpenStudy (jhannybean):

Thanks :)

OpenStudy (anonymous):

haha no problem :) u defiantly earn it